# Demystifying Spectral Feature Learning for Instrumental Variable Regression

### v1

@@ 2.1 Background
[…]
@@ 2.2 On the Conditional Expectation Operator: NPIV as an Ill-posed Inverse Problem
As highlighted by Eq. 2, this operator plays a central role in the NPIV model. It encodes the conditional distribution of $X$ given $Z$ and provides a convenient lens through which to analyze how efficiently one can estimate $h_{0}$. The NPIV equation admits a solution if and only if $r_{0}\in\mathcal{R}(\mathcal{T})$, and is identifiable if and only if $\mathcal{T}$ is injective. We therefore adopt the following assumption.
@@
The existence part is essential: without it, there is no structural function $h_{0}$ satisfying the model in Eq. 1. While injectivity can in principle be relaxed, this comes at the cost of recovering a different object. Specifically, if $h_{0}$ is not identifiable, one may instead recover the minimum-norm solution to Eq. 2, namely $h_{*}\doteq\mathcal{T}^{\dagger}r_{0}$ [14]. [^2] This corresponds to the unique element in the pre-image of $r_{0}$ under $\mathcal{T}$ that minimizes the $L_{2}$-norm and is equivalently characterized as the unique element of the set $\mathcal{T}^{-1}(\{r_{0}\})\cap\mathcal{N}(\mathcal{T})^{\perp}$ [2]. We assume injectivity henceforth to simplify the discussion.
Eq. 2 is an inverse problem in which both the operator $\mathcal{T}$ and the right-hand side $r_{0}$ are unknown. This inverse problem is typically ill-posed, as $\mathcal{T}^{-1}$ is generally not continuous. [^3] In particular, $\mathcal{T}^{-1}$ is discontinuous if $\mathcal{T}$ is compact and not finite rank. Hence, even if $\mathcal{T}$ were known, an approximation $\hat{r}\approx r_{0}$ would not ensure that $\mathcal{T}^{-1}\hat{r}$ is close to $h_{0}=\mathcal{T}^{-1}r_{0}$. In such settings, regularization is required to obtain stable approximate inverses. Two common regularization approaches are Tikhonov regularization, where $\mathcal{T}^{-1}\approx(\mathcal{T}^{\star}\mathcal{T}+\lambda I)^{-1}\mathcal{T}^{\star}$, $\lambda>0$, and spectral cut-off, which truncates the Singular Value Decomposition (SVD) [13].
@@
Assumption 2 is standard in the NPIV literature (Assumption A.1 in [10]). It may fail when $X$ and $Z$ share variables or are deterministically related. For example, if $Z=(W,V)$ and $X=V$, then the joint distribution is not dominated by the product measure.
@@
Under Assumption 2, $\mathcal{T}$ is a Hilbert–Schmidt operator and thus compact.
@@
Compact Decomposition and Eigensubspaces. As $\mathcal{T}$ is a conditional expectation operator, it is non-expansive, i.e., $\|\mathcal{T}\|_{\text{op}}\leq 1$. Under Assumption 2, $\mathcal{T}$ is compact, and therefore admits a SVD
@@
where $\{u_{i}\}\subset L_{2}(Z)$ and $\{v_{i}\}\subset L_{2}(X)$ are orthonormal systems (left and right singular functions, respectively) spanning $\mathcal{N}(\mathcal{T})^{\perp}$ and $\overline{\mathcal{R}(\mathcal{T})}$ respectively, and $\sigma_{i}$ are the nonnegative singular values in nonincreasing order. The leading singular triplet is known explicitly: $\sigma_{1}=1$, $u_{1}=\mathbbm{1}_{\mathcal{Z}}$ and $v_{1}=\mathbbm{1}_{\mathcal{X}}$ corresponding to the constant functions. If $\mathcal{T}$ is infinite-rank, then $\sigma_{i}\to 0$ as $i\to\infty$. Under the assumption that $\mathcal{T}$ is injective (Assumption 1), we have $\mathcal{N}(\mathcal{T})=\{0\}$, hence $\{v_{i}\}_{i\geq 1}$ forms an orthonormal basis (ONB) of $L_{2}(X)$ and $\sigma_{i}>0$ for all $i\geq 1$. Finally, since $\mathcal{T}$ is Hilbert–Schmidt, $\sum_{i\geq 1}\sigma_{i}^{2}<\infty$. For any $d\geq 1$, define the leading eigensubspaces of $\mathcal{T}$ as
@@ 3.1 Sieve 2SLS estimator
where the minimizer, $h_{0}$, is unique under Assumption 1. A popular strategy to estimate $h_{0}$ is to consider a hypothesis space $\mathcal{H}_{\mathcal{X}}\subset L_{2}\left(X\right)$ and a procedure to estimate the action of $\mathcal{T}$: $\hat{\mathcal{T}}h\approx\mathcal{T}h$ for all $h\in\mathcal{H}_{\mathcal{X}}$, in order to obtain
@@
possibly subject to regularization. In the algorithm 2SLS, we consider features on both $\mathcal{X}$ and $\mathcal{Z}$, and both steps (estimating $h\mapsto\mathcal{T}h$ and solving Eq. 4) are carried out via linear regression in feature space. The two sets of features may be infinite-dimensional—e.g., using Reproducing Kernel Hilbert Spaces (RKHS)—or finite-dimensional—e.g., using splines, wavelets, or neural network features. When the feature dimension grows with sample size, we obtain a sieve. Since our focus is on neural network-based features, we present the finite-dimensional version, noting that the ideas naturally extend to infinite-dimensional cases with proper regularization [33, 29]. We refer to Appendix B for a description of other 2SLS methods.
@@
The pseudo-inverse yields the minimum-norm solution in the case of non-uniqueness. Regularized versions, such as Tikhonov-regularized least squares (also known as ridge regression), can be obtained by penalizing with $\eta\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$ for $\eta>0$. Computing $\hat{A}$ is referred to as the first stage of the 2SLS algorithm. For the second stage, for any $h=\theta^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)\in\mathcal{H}_{\mathcal{X}}$, we define $\hat{\mathcal{T}}h(Z)\doteq\theta^{\mathsf{\scriptscriptstyle T}}\hat{F}(Z)$ with $\hat{F}(Z)=\hat{A}\psi(Z)$. Plugging this into Eq. 4 yields $\hat{h}=\hat{\theta}^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)$ with
@@
Albeit looking different, solving this min-max problem in closed-form leads to $\hat{\theta}_{bis}=\hat{\theta}$ (when we introduce regularization in Eq. 5), and thus $\hat{h}_{bis}=\hat{h}$. The saddle-point form could be useful for deriving convergence rates, see [2] for example, as it allows us to use the theory of saddle-point problems. However, it is not necessary for the implementation of the algorithm.
@@ 3.2 Convergence rates for Sieve 2SLS
Assumption 3-(i) is a condition on the approximation properties of the features used for the instrument space. It says the image of $\mathcal{H}_{\varphi,d}$ through $\mathcal{T}$ lies almost entirely inside $\mathcal{H}_{\psi,d}$, i.e., the projection onto the orthogonal complement vanishes faster than the inverse ill-posedness. Assumption 3-(ii) is a stability condition sometimes referred as link condition that is standard in the NPIV literature (see Assumption 6 in [3], Assumption 5.2(ii) in [7] and Assumption 3 in [22]). It says that $\mathcal{T}$ becomes increasingly contractive on the component of $h_{0}$ lying outside the sieve space as the sieve dimension increases. The next assumption ensures the conditional variance of the noise is uniformly bounded.
@@
Let $\hat{h}$ be the 2SLS estimator from Eq. 5, using features $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and $\psi:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose Assumptions 1, 3 and 4 hold and $\tau_{\varphi,d}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$, where $\zeta_{\varphi,\psi,d}\doteq\max\{\sup_{x}\left\|\mathbb{E}[\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\varphi(x)\right\|_{\ell_{2}},\ \sup_{z}\left\|\mathbb{E}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\psi(z)\right\|_{\ell_{2}}\}$. Then:
@@ 3.3 Convergence rates for Sieve 2SLS with spectral features
A drawback of using non-adaptive features (such as cosine or spline bases) is that verifying Assumption 3 or characterizing the growth of $\tau_{\varphi,d}$ is generally difficult. In contrast, we now consider features constructed from the SVD of $\mathcal{T}$, for which both conditions can be exactly characterized. Moreover, while estimators based on universal bases can achieve minimax optimality under smoothness assumptions on $\mathcal{T}$ and $h_{0}$, they are often outperformed in practice by estimators using data-dependent features. This highlights that minimax rates, though important, do not capture the full picture of empirical performance. Recall from Eq. 3 the singular value decomposition of $\mathcal{T}$, and the definition of the top-$d$ right eigenspace $\mathcal{V}_{d}$. The following result shows that this subspace achieves the minimal possible measure of ill-posedness among all $d$-dimensional subspaces of $L_{2}(X)$.
@@
Let Assumption 2 hold. For any $d\geq 1$, the smallest possible value of $\tau_{\varphi,d}$ is $\tau_{\varphi,d}=\sigma_{d}^{-1}$, achieved when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$.
@@
Let $\hat{h}$ be the 2SLS estimator from Eq. 5 using spectral features. Let Assumptions 1, 2 and 4 hold and $\sigma_{d}^{-1}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$. Then:
@@ 4 Sieve 2SLS with learned spectral features
Let Assumption 2 hold and let $d\geq 1$ be such that $\sigma_{d}>\sigma_{d+1}$. For all feature maps $\varphi:\mathcal{X}\to\mathbb{R}^{d}$, $\psi:\mathcal{Z}\to\mathbb{R}^{d}$, we have
@@
where equality holds if and only if $\mathcal{T}_{d}(\varphi,\psi)=\mathcal{T}_{d}$, in which case $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$, $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$. $\mathcal{T}_{d}$ is the rank-$d$ truncation of $\mathcal{T}$ introduced in Section 2.
While the Eckart–Young–Mirsky objective in Eq. 6 offers a clean objective to learn spectral features, it appears impractical due to the unknown nature of the operator $\mathcal{T}$. However, this objective admits an equivalent formulation as a spectral contrastive learning loss [35]:
@@
^$\hat{\mathcal{L}}_{d}(\varphi,\psi)\doteq\frac{1}{m(m-1)}\sum_{i\neq j}\left(\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{j})\right)^{2}-\frac{2}{m}\sum_{i=1}^{m}\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{i}).$ (8)
The spectral contrastive loss is the basis of the recent algorithm proposed in [34], where $\varphi$ and $\psi$ are parametrized by neural networks and optimized via stochastic gradient descent. While [34] motivate this objective heuristically, Eq. 6 shows that it directly targets the leading singular structure of $\mathcal{T}$. The use of objectives similar to Eq. 6 for operator learning or Eq. 7 for representation learning has a rich history that we detail in Appendix B.
Comparison to [34]. To justify the use of the contrastive loss in Eq. 8, [34] assume that the density $p(x,z)$ factorizes as
@@
We argue that this assumption is overly strong: it holds if and only if $\mathcal{T}$ has rank at most $d$, in which case the spectral features perfectly capture the structure of $\mathcal{T}$. This result is formalized in 5, Appendix A.3. We should instead interpret Eq. (9) as an approximation rather than a literal assumption. In practice, the learned features $\varphi,\psi$ provide a rank-$d$ approximation of $\mathcal{T}$, and the quality of this approximation governs the performance of the resulting Sieve 2SLS estimator.
Let $\hat{\varphi},\hat{\psi}$ be $d$-dimensional features trained with loss Eq. 8 with $\tilde{D}_{m}$ and define $\hat{\mathcal{T}}_{d}\doteq\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. We make the following assumption:
@@
Let Assumptions 1, 2, 4 and 5 hold and let $\hat{\varepsilon}_{d}\doteq\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}$ be such that $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$.
@@
    Let $\hat{h}$ be the 2SLS estimator from Eq. 5, using features $\hat{\varphi}:\mathcal{X}\to\mathbb{R}^{d}$ and $\hat{\psi}:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose $(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}\zeta_{\hat{\varphi},\hat{\psi},d}\sqrt{(\log d)/n}=o(1)$ and $\hat{\varepsilon}_{d}=o_{d}(\sigma_{d}^{2})$, then
@@
We postpone the proof to Appendix A.3. The result quantifies how close the learned features must be to $\mathcal{T}_{d}$ in order to preserve the desirable properties of spectral 2SLS. Theorem 3-i) shows that as long as the learned operator $\hat{\mathcal{T}}_{d}$ is a good approximation of the true rank-$d$ truncation $\mathcal{T}_{d}$, the learned features inherit a favorable measure of ill-posedness. Theorem 3-ii) further ensures that if contrastive learning yields a near-optimal rank-$d$ approximation of $\mathcal{T}$, the downstream Sieve 2SLS estimator retains strong statistical guarantees. Importantly, the spectral features generalization error $\hat{\varepsilon}_{d}$ encapsulates the statistical and optimization errors in learning spectral features. It depends on the number of samples used to train the contrastive loss, the expressivity and architecture of the neural networks parametrizing $\varphi$ and $\psi$, and the complexity of the leading eigensubspaces of $\mathcal{T}$. As $\mathcal{L}_{d}(\varphi,\psi)$ converges to $\mathcal{L}_{d}^{\mathrm{min}}\doteq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ $\hat{\varepsilon}_{d}$ converges to 00 by Theorem 2. Precisely controlling $\hat{\varepsilon}_{d}$ remains an important question for future work.

### v2

@@ 2.1 Background
Probability Spaces and Function Spaces. $Y$ is defined on $\mathbb{R}$, while $X$ and $Z$ take values in measurable spaces $\mathcal{X}$ and $\mathcal{Z}$, respectively, endowed with their $\sigma$-fields $\mathcal{F}_{\mathcal{X}}$ and $\mathcal{F}_{\mathcal{Z}}$. We let $(\Omega,\mathcal{F},\mathbb{P})$ be the underlying probability space with expectation operator $\mathbb{E}$. For $R,S\in\{X,Z,(X,Z)\}$, let $\pi_{R}$ denote the push-forward of $\mathbb{P}$ under $R$, and let $\pi_{R}\otimes\pi_{S}$ denote the product measure on the corresponding product $\sigma$-field, defined by $(\pi_{R}\otimes\pi_{S})(A\times B)=\pi_{R}(A)\pi_{S}(B)$ for measurable sets $A,B$. For $R\in\{X,Z\}$ and corresponding domain $\mathcal{R}\in\left\{\mathcal{X},\mathcal{Z}\right\}$, we abbreviate $L_{2}(R)\equiv L_{2}(\mathcal{R},\pi_{R})$ as the space of measurable functions $f:\mathcal{R}\to\mathbb{R}$ such that $\int|f|^{2}\,d\pi_{R}<\infty$, defined up to $\pi_{R}$-almost everywhere equivalence.
@@ 2.2 On the Conditional Expectation Operator: NPIV as an Ill-posed Inverse Problem
As highlighted by Eq.˜2, this operator plays a central role in the NPIV model. It encodes the conditional distribution of $X$ given $Z$ and provides a convenient lens through which to analyze how efficiently one can estimate $h_{0}$. The NPIV equation admits a solution if and only if $r_{0}\in\mathcal{R}(\mathcal{T})$, and is identifiable if and only if $\mathcal{T}$ is injective. We therefore adopt the following assumption.
@@
The existence part is essential: without it, there is no structural function $h_{0}$ satisfying the model in Eq.˜1. While injectivity can in principle be relaxed, this comes at the cost of recovering a different object. Specifically, if $h_{0}$ is not identifiable, one may instead recover the minimum-norm solution to Eq.˜2, namely $h_{*}\doteq\mathcal{T}^{\dagger}r_{0}$ [14]. [^2] This corresponds to the unique element in the pre-image of $r_{0}$ under $\mathcal{T}$ that minimizes the $L_{2}$-norm and is equivalently characterized as the unique element of the set $\mathcal{T}^{-1}(\{r_{0}\})\cap\mathcal{N}(\mathcal{T})^{\perp}$ [2]. We assume injectivity henceforth to simplify the discussion.
Eq.˜2 is an inverse problem in which both the operator $\mathcal{T}$ and the right-hand side $r_{0}$ are unknown. This inverse problem is typically ill-posed, as $\mathcal{T}^{-1}$ is generally not continuous. [^3] In particular, $\mathcal{T}^{-1}$ is discontinuous if $\mathcal{T}$ is compact and not finite rank. Hence, even if $\mathcal{T}$ were known, an approximation $\hat{r}\approx r_{0}$ would not ensure that $\mathcal{T}^{-1}\hat{r}$ is close to $h_{0}=\mathcal{T}^{-1}r_{0}$. In such settings, regularization is required to obtain stable approximate inverses. Two common regularization approaches are Tikhonov regularization, where $\mathcal{T}^{-1}\approx(\mathcal{T}^{\star}\mathcal{T}+\lambda I)^{-1}\mathcal{T}^{\star}$, $\lambda>0$, and spectral cut-off, which truncates the Singular Value Decomposition (SVD) [13].
@@
Assumption˜2 is standard in the NPIV literature (Assumption A.1 in [10]). It may fail when $X$ and $Z$ share variables or are deterministically related. For example, if $Z=(W,V)$ and $X=V$, then the joint distribution is not dominated by the product measure.
@@
Under Assumption˜2, $\mathcal{T}$ is a Hilbert–Schmidt operator and thus compact.
@@
Compact Decomposition and Eigensubspaces. As $\mathcal{T}$ is a conditional expectation operator, it is non-expansive, i.e., $\|\mathcal{T}\|_{\text{op}}\leq 1$. Under Assumption˜2, $\mathcal{T}$ is compact, and therefore admits a SVD
@@
where $\{u_{i}\}\subset L_{2}(Z)$ and $\{v_{i}\}\subset L_{2}(X)$ are orthonormal systems (left and right singular functions, respectively) spanning $\mathcal{N}(\mathcal{T})^{\perp}$ and $\overline{\mathcal{R}(\mathcal{T})}$ respectively, and $\sigma_{i}$ are the nonnegative singular values in nonincreasing order. The leading singular triplet is known explicitly: $\sigma_{1}=1$, $u_{1}=\mathbbm{1}_{\mathcal{Z}}$ and $v_{1}=\mathbbm{1}_{\mathcal{X}}$ corresponding to the constant functions. If $\mathcal{T}$ is infinite-rank, then $\sigma_{i}\to 0$ as $i\to\infty$. Under the assumption that $\mathcal{T}$ is injective (Assumption˜1), we have $\mathcal{N}(\mathcal{T})=\{0\}$, hence $\{v_{i}\}_{i\geq 1}$ forms an orthonormal basis (ONB) of $L_{2}(X)$ and $\sigma_{i}>0$ for all $i\geq 1$. Finally, since $\mathcal{T}$ is Hilbert–Schmidt, $\sum_{i\geq 1}\sigma_{i}^{2}<\infty$. For any $d\geq 1$, define the leading eigensubspaces of $\mathcal{T}$ as
@@ 3.1 Sieve 2SLS estimator
where the minimizer, $h_{0}$, is unique under Assumption˜1. A popular strategy to estimate $h_{0}$ is to consider a hypothesis space $\mathcal{H}_{\mathcal{X}}\subset L_{2}\left(X\right)$ and a procedure to estimate the action of $\mathcal{T}$: $\hat{\mathcal{T}}h\approx\mathcal{T}h$ for all $h\in\mathcal{H}_{\mathcal{X}}$, in order to obtain
@@
possibly subject to regularization. In the algorithm 2SLS, we consider features on both $\mathcal{X}$ and $\mathcal{Z}$, and both steps (estimating $h\mapsto\mathcal{T}h$ and solving Eq.˜4) are carried out via linear regression in feature space. The two sets of features may be infinite-dimensional—e.g., using Reproducing Kernel Hilbert Spaces (RKHS)—or finite-dimensional—e.g., using splines, wavelets, or neural network features. When the feature dimension grows with sample size, we obtain a sieve. Since our focus is on neural network-based features, we present the finite-dimensional version, noting that the ideas naturally extend to infinite-dimensional cases with proper regularization [33, 29]. We refer to Appendix B for a description of other 2SLS methods.
@@
The pseudo-inverse yields the minimum-norm solution in the case of non-uniqueness. Regularized versions, such as Tikhonov-regularized least squares (also known as ridge regression), can be obtained by penalizing with $\eta\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}$ for $\eta>0$. Computing $\hat{A}$ is referred to as the first stage of the 2SLS algorithm. For the second stage, for any $h=\theta^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)\in\mathcal{H}_{\mathcal{X}}$, we define $\hat{\mathcal{T}}h(Z)\doteq\theta^{\mathsf{\scriptscriptstyle T}}\hat{F}(Z)$ with $\hat{F}(Z)=\hat{A}\psi(Z)$. Plugging this into Eq.˜4 yields $\hat{h}=\hat{\theta}^{\mathsf{\scriptscriptstyle T}}\varphi(\cdot)$ with
@@
Albeit looking different, solving this min-max problem in closed-form leads to $\hat{\theta}_{bis}=\hat{\theta}$ (when we introduce regularization in Eq.˜5), and thus $\hat{h}_{bis}=\hat{h}$. The saddle-point form could be useful for deriving convergence rates, see [2] for example, as it allows us to use the theory of saddle-point problems. However, it is not necessary for the implementation of the algorithm.
@@ 3.2 Convergence rates for Sieve 2SLS
Assumption˜3-(i) is a condition on the approximation properties of the features used for the instrument space. It says the image of $\mathcal{H}_{\varphi,d}$ through $\mathcal{T}$ lies almost entirely inside $\mathcal{H}_{\psi,d}$, i.e., the projection onto the orthogonal complement vanishes faster than the inverse ill-posedness. Assumption˜3-(ii) is a stability condition sometimes referred as link condition that is standard in the NPIV literature (see Assumption 6 in [3], Assumption 5.2(ii) in [7] and Assumption 3 in [22]). It says that $\mathcal{T}$ becomes increasingly contractive on the component of $h_{0}$ lying outside the sieve space as the sieve dimension increases. The next assumption ensures the conditional variance of the noise is uniformly bounded.
@@
Let $\hat{h}$ be the 2SLS estimator from Eq.˜5, using features $\varphi:\mathcal{X}\to\mathbb{R}^{d}$ and $\psi:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose Assumptions˜1, 3 and 4 hold and $\tau_{\varphi,d}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$, where $\zeta_{\varphi,\psi,d}\doteq\max\{\sup_{x}\left\|\mathbb{E}[\varphi(X)\varphi(X)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\varphi(x)\right\|_{\ell_{2}},\ \sup_{z}\left\|\mathbb{E}[\psi(Z)\psi(Z)^{\mathsf{\scriptscriptstyle T}}]^{-1/2}\psi(z)\right\|_{\ell_{2}}\}$. Then:
@@ 3.3 Convergence rates for Sieve 2SLS with spectral features
A drawback of using non-adaptive features (such as cosine or spline bases) is that verifying Assumption 3 or characterizing the growth of $\tau_{\varphi,d}$ is generally difficult. In contrast, we now consider features constructed from the SVD of $\mathcal{T}$, for which both conditions can be exactly characterized. Moreover, while estimators based on universal bases can achieve minimax optimality under smoothness assumptions on $\mathcal{T}$ and $h_{0}$, they are often outperformed in practice by estimators using data-dependent features. This highlights that minimax rates, though important, do not capture the full picture of empirical performance. Recall from Eq.˜3 the singular value decomposition of $\mathcal{T}$, and the definition of the top-$d$ right eigenspace $\mathcal{V}_{d}$. The following result shows that this subspace achieves the minimal possible measure of ill-posedness among all $d$-dimensional subspaces of $L_{2}(X)$.
@@
Let Assumption˜2 hold. For any $d\geq 1$, the smallest possible value of $\tau_{\varphi,d}$ is $\tau_{\varphi,d}=\sigma_{d}^{-1}$, achieved when $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$.
@@
Let $\hat{h}$ be the 2SLS estimator from Eq.˜5 using spectral features. Let Assumptions˜1, 2 and 4 hold and $\sigma_{d}^{-1}\zeta_{\varphi,\psi,d}\sqrt{(\log d)/n}=o(1)$. Then:
@@ 4 Sieve 2SLS with learned spectral features
Let Assumption˜2 hold and let $d\geq 1$ be such that $\sigma_{d}>\sigma_{d+1}$. For all feature maps $\varphi:\mathcal{X}\to\mathbb{R}^{d}$, $\psi:\mathcal{Z}\to\mathbb{R}^{d}$, we have
@@
where equality holds if and only if $\mathcal{T}_{d}(\varphi,\psi)=\mathcal{T}_{d}$, in which case $\mathcal{H}_{\varphi,d}=\mathcal{V}_{d}$, $\mathcal{H}_{\psi,d}=\mathcal{U}_{d}$. $\mathcal{T}_{d}$ is the rank-$d$ truncation of $\mathcal{T}$ introduced in Section˜2.
While the Eckart–Young–Mirsky objective in Eq.˜6 offers a clean objective to learn spectral features, it appears impractical due to the unknown nature of the operator $\mathcal{T}$. However, this objective admits an equivalent formulation as a spectral contrastive learning loss [35]:
@@
$\hat{\mathcal{L}}_{d}(\varphi,\psi)\doteq\frac{1}{m(m-1)}\sum_{i\neq j}\left(\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{j})\right)^{2}-\frac{2}{m}\sum_{i=1}^{m}\varphi(\tilde{x}_{i})^{\mathsf{\scriptscriptstyle T}}\psi(\tilde{z}_{i}).$ (8)
The spectral contrastive loss is the basis of the recent algorithm proposed in [34], where $\varphi$ and $\psi$ are parametrized by neural networks and optimized via stochastic gradient descent. While [34] motivate this objective heuristically, Eq.˜6 shows that it directly targets the leading singular structure of $\mathcal{T}$. The use of objectives similar to Eq.˜6 for operator learning or Eq.˜7 for representation learning has a rich history that we detail in Appendix B.
Comparison to [34]. To justify the use of the contrastive loss in Eq.˜8, [34] assume that the density $p(x,z)$ factorizes as
@@
We argue that this assumption is overly strong: it holds if and only if $\mathcal{T}$ has rank at most $d$, in which case the spectral features perfectly capture the structure of $\mathcal{T}$. This result is formalized in ˜5, Appendix A.3. We should instead interpret Eq. (9) as an approximation rather than a literal assumption. In practice, the learned features $\varphi,\psi$ provide a rank-$d$ approximation of $\mathcal{T}$, and the quality of this approximation governs the performance of the resulting Sieve 2SLS estimator.
Let $\hat{\varphi},\hat{\psi}$ be $d$-dimensional features trained with loss Eq.˜8 with $\tilde{D}_{m}$ and define $\hat{\mathcal{T}}_{d}\doteq\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. We make the following assumption:
@@
Let Assumptions˜1, 2, 4 and 5 hold and let $\hat{\varepsilon}_{d}\doteq\left\|\hat{\mathcal{T}}_{d}-\mathcal{T}_{d}\right\|_{\mathrm{op}}$ be such that $\hat{\varepsilon}_{d}<(1-1/\sqrt{2})\sigma_{d}$.
@@
    Let $\hat{h}$ be the 2SLS estimator from Eq.˜5, using features $\hat{\varphi}:\mathcal{X}\to\mathbb{R}^{d}$ and $\hat{\psi}:\mathcal{Z}\to\mathbb{R}^{d}$. Suppose $(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}\zeta_{\hat{\varphi},\hat{\psi},d}\sqrt{(\log d)/n}=o(1)$ and $\hat{\varepsilon}_{d}=o_{d}(\sigma_{d}^{2})$, then
@@
We postpone the proof to Appendix A.3. The result quantifies how close the learned features must be to $\mathcal{T}_{d}$ in order to preserve the desirable properties of spectral 2SLS. Theorem˜3-i) shows that as long as the learned operator $\hat{\mathcal{T}}_{d}$ is a good approximation of the true rank-$d$ truncation $\mathcal{T}_{d}$, the learned features inherit a favorable measure of ill-posedness. Theorem˜3-ii) further ensures that if contrastive learning yields a near-optimal rank-$d$ approximation of $\mathcal{T}$, the downstream Sieve 2SLS estimator retains strong statistical guarantees. Importantly, the spectral features generalization error $\hat{\varepsilon}_{d}$ encapsulates the statistical and optimization errors in learning spectral features. It depends on the number of samples used to train the contrastive loss, the expressivity and architecture of the neural networks parametrizing $\varphi$ and $\psi$, and the complexity of the leading eigensubspaces of $\mathcal{T}$. As $\mathcal{L}_{d}(\varphi,\psi)$ converges to $\mathcal{L}_{d}^{\mathrm{min}}\doteq\left\|\mathcal{T}-\mathcal{T}_{d}\right\|_{\scriptscriptstyle\mathrm{HS}}^{2},$ $\hat{\varepsilon}_{d}$ converges to 0 by Theorem˜2. Precisely controlling $\hat{\varepsilon}_{d}$ remains an important question for future work.